OpenStudy (anmolsingh):

If the displacement x=(1/2)vt, where v is the velocity and t is the time, Then proove that the acceleration is constant

OpenStudy (anonymous):

x=(1/2)vt diff wrt t dx/dt=(1/2) v (dt/dt)+(1/2) (dv/dt) t v=(1/2)v+(1/2)at v=at diff wrt t (dv/dt) = (da/dt) t+a (dt/dt) a = (da/dt) t + a ==>da/dt = 0 ==>change in a is zero ==> a is constant

OpenStudy (anmolsingh):

THANKS!!!